Suppose that you hold a piece of land in the City of London that you may want to sell in one year. As a U.S. resident, you are concerned with the dollar value of the land. Assume that, if the British economy booms in the future, the land will be worth £1,824 and one British pound will be worth $1.41. If the British economy slows down, on the other hand, the land will be worth less, i.e., £1,548, but the pound will be stronger, i.e., $1.62/£. You feel that the British economy will experience a boom with a 60% probability and a slow-down with a 40% probability.

(b) Compute the variance of the dollar value of your property that is attributable to the exchange rate uncertainty.
Answer #1

Total Value of land = Value * Pound

For Boom = 1824 * 1.41 = 2571.84

For Slow down = 1548 * 1.62 = 2507.76

E(P) = Probability * Value of land

= 60%*2571.84 + 40%*2507.76

= 1543.10 + 1003.10

= 2546.20

E(S) = Probability * Value of pound

= 60%*1.41 + 40%*1.62

= 0.846+0.648

= 1.494

Var(S) = Probability [Bond Value - E(S)]2

= 60%(1.41-1.49)2 + 40%(1.62-1.49)2

= 0.00384 + 0.00676

= 0.0106

Cov(P,S) = Probability*(Value of land - E(P))(Value of bond - E(S))

= 60%(2571.84 - 2546.20)(1.41-1.49) + 40%(2507.76 - 2546.20)(1.62-1.49)

= 60%*25.64*-0.08 + 40% * -38.44*0.13

= -1.23 - 2.00

= -3.23

b = Cov(P,S) / Var(S)

= -3.23 / 0.0106

= 304.72

Variance = b2 * Var(S)

= (304.72)2 * 0.0106

= 984.26